Selected data from Franklin Company follow: Balance Sheets As of December 31 2018 2017 Accounts receivable Allowance for doubtful accounts Net accounts receivable Inventories, lower of cost or market $404,000 $379,000 (20,200)(15,160) $383,800 $363,840 $476,000 $435,000 Income Statement For the Years Ended December 31 2018 2017 Net credit sales Net cash sales Net sales $2,007,000 $1,752,000 415,000 309,00 Cost of goods sold Selling, general, and administrative expenses Other expenses 2,422,000 1,606,000 2,061,000 1,432,000 214,600 239,900 40,80022,500 Total operating expenses $1,886,700 $1,669,100
media%2F546%2F54699af2-80dd-4ec4-85e5-a5
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Accounts receivables, net:

A

2017

$          363,840.00

B

2018

$          383,800.00

C = A+B

Total

$          747,640.00

D = C/2

Average Accounts receivables

$          373,820.00

E

Net Credit Sale 2018

$      2,007,000.00

F = E/D

Accounts receivables Turnover ratio

5.37 times = answer

  • Requirement ‘b’

Inventories balance:

A

2017

$          435,000.00

B

2018

$          476,000.00

C = A+B

Total

$          911,000.00

D = C/2

Average Inventories

$          455,500.00

E

Cost of Goods Sold for 2018

$      1,606,000.00

F = E/D

Inventory Turnover Ratio

3.53 times = answer

  • Requirement ‘c’

Working

2017

A

Total Net Sales

$      2,061,000.00

B

Total Operating Expenses

$      1,669,100.00

C = A/B

Net Income

$          391,900.00

D = (C/A) x 100

Net Margin for 2017

19.02%